Mary is now 6 times as old as Nancy was eight years ago. How old is Nancy today if Mary is 30 years old?
Answer by ikleyn(52803)

.
From the condition, you have this equation


    M = 6*(N-8),


and you know that M = 30.


Hence, the equation becomes


    30 = 6*(N-8),


and you can easily solve it


    30/6 = N - 8


    5    = N - 8


    N = 5 + 8 = 13.


ANSWER.  Nancy is 13 years old.